number of H - ions in 5.44 g SrH2



1
Expert's answer
2021-10-29T02:20:45-0400

The molar mass of SrH2 is 89.64 g/mol, and there are 2 moles of H- ions per 1 mole of SrH2. Therefore,

"Number\ of\ H^-=5.44\ g(SrH_2)\times\frac{1\ mol(SrH_2)}{89.64\ g(SrH_2)}\times\frac{2\ mol(H^-)}{1\ mol(SrH_2)}\times\frac{6.022\times10^{23}\ ions(H^-)}{1\ mol(H^-)}=7.31\times10^{22}\ ions"


Answer: "7.31\times10^{22}"